OpenStudy (anonymous):

Find an equation of the line passing through the given points. Use function notation to write the equation. (-2,-4) and (-4,-3)

hero (hero):

y = mx + b -3 = -4m + b -4 = -2m + b b = 4m - 3 b = 2m - 4 4m - 3 = 2m - 4 4m - 2m = 3 - 4 2m = -1 m = -1/2 b = 4(-1/2) - 3 b = 2(-1) = -2 y = -1/2(x) - 2